40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16     Entry time: Thu Oct 25 23:35:36 2007
Author: waldman 
Type: Other 
Category: OMC 
Subject: Hang the OMC! 
[Pinkesh, Sam]

We tried, convicted and hung the OMC today. The OMC was found guilty of being overweight, and unsymmetrically balanced. The unsymmetry was kind of expected and was corrected with a hefty stack of counterweights positioned over the counterweighting holes. The stacks will be measured at some future date and correctly sized objects machined. The overweightness showed up when the level hanging breadboard was about 5 mm low. This showed up in the board height above the table as well as the OSEM flag positions within their holes. The problem was remedied with a liposuction of the intermediate mass. We removed both small vertical cylinder weights that Chris added, and then we removed the heavy steel transverse weight that can be used to adjust the tip around the long axis (I forgot what its called).

The top of the breadboard ended up about 154 mm off the table. The breadboard is 39 mm thick, and the optics are centered (30 - 12.7) = 17.3 mm below the surface for a as hanging beams height of 154 -39 - 17.3 = 97.7 mm or about an 0.150 inches lower than we were aiming for. Can I get a refund?

We screwed up in multiple ways:
  • The slotted disks that capture the wires do not have the alignment bore used to center the wire in the hole
  • We didn't correctly route the far field QPD cable so it runs funny
  • We didn't have a tool which could be used to get two of the DCPD preamp box mounting screws (which are M3's chub!)
  • We don't have the cable clamps to tie off the electrical cables to the intermediate mass
  • We don't have any of the cabling from the OMC-SUS top to the rack so we can't test anything
  • We haven't uploaded pretty pictures for all to see

We left the OMC partially suspended by the OMC-SUS and partly resting on the installation lab jacks which are currently acting as EQ stops. After we fix the cabling we will more permanently hang it. PS, It looks like the REFL beam extraction will be tricky so we need to get on that....
Attachment 1: IMG_1483.jpg  562 kB  Uploaded Fri Oct 26 01:00:30 2007  | Hide | Hide all
IMG_1483.jpg
Attachment 2: IMG_1481.jpg  555 kB  Uploaded Fri Oct 26 01:01:19 2007  | Hide | Hide all
IMG_1481.jpg
ELOG V3.1.3-